OpenStudy (anonymous):

write the equation of the line in slope intercept form that is perpendicular to the line y=3/5x-1 and passes through the point (-9,4)

OpenStudy (anonymous):

is it (3/5)x

OpenStudy (anonymous):

so would it be y=3/5x+47/5

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

When it states perpendicular, the slope would be the inverse negative.

OpenStudy (anonymous):

So for the slope of the perpendicular line, the slope would be -5/3 x

OpenStudy (anonymous):

so then would it be y=-5/3x-11

OpenStudy (anonymous):

or y=-5/3x+19

OpenStudy (anonymous):

I don't know. I would have to plug it in to find out.

OpenStudy (anonymous):

But the slope is correct.

OpenStudy (anonymous):

would you please do that:)

OpenStudy (anonymous):

I would assume its -5/3x-11, because the slope is negative, and it would never cross the x axis at 19.

OpenStudy (anonymous):

alright:)

OpenStudy (anonymous):

Just this confirm, this is a straight line, correct?

OpenStudy (anonymous):

I believe so